### v1

# Towards General Industrial Intelligence: A Survey on Industrial IoT-Enhanced Continual Large Models

Abstract
Currently, most applications in the Industrial Internet of Things (IIoT) still rely on CNN-based neural networks. Although Transformer-based large models (LMs), including language, vision, and multimodal models, have demonstrated impressive capabilities in AI-generated content (AIGC), their application in industrial domains, such as detection, planning, and control, remains relatively limited. Deploying pre-trained LMs in industrial environments often encounters the challenge of stability and plasticity due to the complexity of tasks, the diversity of data, and the dynamic nature of user demands. To address these challenges, the pre-training and fine-tuning strategy, coupled with continual learning, has proven to be an effective solution, enabling models to adapt to dynamic demands while continuously optimizing their inference and decision-making capabilities. This paper surveys the integration of LMs into IIoT-enhanced General Industrial Intelligence (GII), focusing on two key areas: LMs for GII and LMs on GII. The former focuses on leveraging LMs to provide optimized solutions for industrial application challenges, while the latter investigates continuous optimization of LMs learning and inference capabilities in collaborative scenarios involving industrial devices, edge computing, and cloud computing. This paper provides insights into the future development of GII, aiming to establish a comprehensive theoretical framework and research direction for GII, thereby advancing GII towards a more general and adaptive future.
Index Terms:
General Industrial Intelligence, Large Language Models, Continual Learning, Industrial Internet of Things, Edge Computing.
I Introduction
Industrial 5.0 integrates cyber-physical-social elements into manufacturing, emphasizing the interaction between digital and physical systems as well as human-machine collaboration, effectively connecting devices, objects, and people through the Internet [1]. With the rapid development of the Internet of Things (IIoT) [2, 3, 4], communication technologies [5, 6], AI-generated content (AIGC) [7], robotics and digtal twins [8, 9, 10], modern industrial systems are becoming increasingly complex. These systems not only generate high-frequency unimodal data but also encompass multimodal data types such as text, images, videos, code, and audio. Industrial big data can be leveraged to create digital manufacturing workflows and industrial processes, significantly driving improvements in productivity, efficiency, and effectiveness within Industry 5.0 and cyber-physical-social systems.
Figure 1: Layering the general industrial intelligence ecosystem.
As illustrated in Fig. 1, datasets and models constitute the foundational elements of the GII ecosystem, driving the rapid development of higher-level algorithms and applications. These applications, which include intelligent control systems, predictive maintenance [11], fault diagnosis [12, 13], and anomaly detection [14], rely heavily on the extraction and analysis of IIoT data. The success of GII is particularly dependent on its ability to efficiently learn and extract valuable features from these IIoT datasets.
Large models (LMs) based on the Transformer [15], such as Large Language Models (LLMs) [16, 17, 18], vision models [19, 20], time series models [21], and multimodal models [22, 23], have gained widespread attention due to their unique advantages. With billions to trillions of parameters, these models accumulate extensive knowledge through pre-training on large datasets, significantly advancing the automation and diversification of data processing while reducing reliance on human expertise.
In the industrial sector, the precision and scalability of LMs make them highly effective for enhancing the accuracy of industrial processes. However, deploying pre-trained LMs in industrial settings requires careful adaptation to specific task architectures, dynamic data distributions, and user preferences. Balancing stability and adaptability in these environments remains a significant challenge, despite the advantages of LMs in multitask generalization, few-shot learning, and inference.
Inspired by the success of LMs in natural language processing (NLP), the industrial community has increasingly explored their potential for GII. One approach involves building industry-specific foundational models from scratch [24], though the limited scale of domain-specific data often hinders the development of emergent capabilities. Alternatively, pre-training on large datasets followed by task-specific fine-tuning has shown immense potential for building robust industrial models, significantly improving performance across diverse tasks. This method effectively addresses the challenges posed by data scarcity in specific domains, while accelerating the development of advanced capabilities in industrial applications.
Adapting LMs for industrial tasks represents a significant research direction [25]. The advantages of these models in cross-task generalization, zero-shot/few-shot learning, and inferencing capabilities offer new avenues for addressing challenges in knowledge transfer, data scarcity, and explainability in GII. Continual Large Models (CLMs) play a crucial role in maintaining and evolving the capabilities of these industrial models over time. Pre-trained on extensive datasets and powered by Transformer architectures, CLMs are designed to continuously learn and adapt, ensuring that industrial large models remain relevant and effective in meeting the evolving demands of GII.
I-A Goals of This Survey
This paper aims to establish a comprehensive perspective and conduct an in-depth analysis of IIoT-Enhanced GII. It proposes a bifurcation of GII into two main categories:
$\bullet$ Large Models for General Industrial Intelligence (LMs for GII): This direction focuses on utilizing the advanced data processing and analysis capabilities of LMs to solve optimization problems inherent in industrial applications. Specifically, LMs enhance the intelligence and operational efficiency of IIoT-driven industrial systems by leveraging their ability to process real-time multimodal IIoT data, perform intricate feature extraction, and ensure precise pattern recognition and result validation, ultimately improving measurement accuracy and system performance across diverse industrial settings.
$\bullet$ Large Models on General Industrial Intelligence (LMs on GII): This perspective examines how industrial applications can extend and refine the capabilities of LMs through continuous model operation in collaborative IIoT device-edge-cloud environments. By adopting continual learning (CL) and online learning strategies, models can adapt to new data and environmental changes without the need for costly retraining. This approach conserves computational resources, minimizes latency, and efficiently handles distribution shifts and performance degradation, ensuring robust model performance in dynamic industrial scenarios.
The introduction of these categories is further clarified with an illustration (Fig. 2), which helps delineate the structural differences and operational mechanisms between the two approaches.
Figure 2: A high-level overview of LMs for GII and LMs on GII. The vertical dimension primarily focuses on leveraging the general knowledge of LMs to enhance General Industrial Intelligence. The horizontal dimension concentrates on the continual updating of LMs in long sequence industrial tasks.
Large Models (LMs) with General Industrial Intelligence (GII)LMs for GII (§$\S$§III)Preliminary (§$\S$§II)LMs on GII (§$\S$§IV)Outlook and Future Opportunities (§$\S$§V)Research Roadmap (§$\S$§II-A)Industrial LMs (§$\S$§II-B)Key Tasks of GII (§$\S$§II-C) Classification (§$\S$§II-C1), Detection/Localization (§$\S$§II-C2), Planning (§$\S$§II-C3), General Q&A (§$\S$§II-C4) LMs as a Data Enhancer (§$\S$§III-A)LMs as a Tuning Predictor (§$\S$§III-B)LMs as a Non-Tuning Predictor (§$\S$§III-C)Classical Continual Learning (§$\S$§IV-A)Pre-trained Models Based CL (§$\S$§IV-B)Efficient Inference Algorithm (§$\S$§IV-C)Benchmarks (§$\S$§IV-D) Model Architectures (§$\S$§II-B1), LLMs (§$\S$§II-B2), Vision-based Models (§$\S$§II-B3), Time Series-based Models (§$\S$§II-B4), Multimodal Models (§$\S$§II-B5) Data Denoise (§$\S$§III-A1), Data Selection (§$\S$§III-A2), Data Augmentation (§$\S$§III-A3), Data Generation (§$\S$§III-A4), Reprogramming (§$\S$§III-A5) Alignment Fine-tuning (§$\S$§III-B1), Efficient Tuning (§$\S$§III-B2) Regularization-based Methods (§$\S$§IV-A1), Replay-based Methods (§$\S$§IV-A2), Parameter Isolation Methods (§$\S$§IV-A3) Mixture of Experts (§$\S$§IV-B1), Analytic Learning-based Methods (§$\S$§IV-B2), Prompt-based Methods (§$\S$§IV-B3), Representation-based Methods (§$\S$§IV-B4) Test Time Adaptation (§$\S$§IV-C1), Edge Cloud Collaboration (§$\S$§IV-C2), Federated Learning (§$\S$§IV-C3) Data Generation (§$\S$§V-A), Model/Knowledge Reuse (§$\S$§V-B), Beyond Single Modality (§$\S$§V-C), Learning with Restricted Resources (§$\S$§V-D)
Figure 3: Organization of our survey.
I-B Distinguishing Features of This Survey
CL has gained significant traction as a research topic in recent years, with numerous studies exploring its application across various domains such as equipment health management [26], robotics [27], and streaming data [28]. In the context of LMs, CL has been recognized as crucial due to the impracticality of frequent retraining caused by the high costs associated with these models’ massive scale [29].
Despite the extensive literature on CL, our survey uniquely addresses the continual adaptation of LMs for IIoT-enhanced industrial systems—a field not thoroughly covered by existing literature. This survey is the pioneer in providing a comprehensive and systematic review of CL methods tailored for LMs across four distinct IIoT-enabled industrial scenarios.
TABLE I: Overview of General Industrial Intelligence Perspectives in Different Surveys (based on the hierarchical structure shown in Fig. 1).
Layer Perspectives Surveys
Application Predictive Maintenance Hurtado et al. 2023 [26]
Application Industrial Detection Li et al. 2024 [30]
Application Autonomous Driving Katare et al. 2023 [31]
Application AIGC Xu et al. 2024 [7]
Algorithm Continual Learning Ours
Foundation Model Design Liu et al. 2022 [32]
Foundation Federated Learning Nguyen et al. 2021 [33]
Infrastructure End-Edge-Cloud Wang et al. 2024 [34]
Infrastructure Green AI Mao et al. 2023 [35]
Else Safety Perez et al. 2023 [36]
Outlined in Table I, this survey distinguishes itself through several key contributions:
$\bullet$ Novel Taxonomy: We introduce a novel theoretical framework for GII. By categorizing the application of large models into two dimensions—“LMs for GII” and “LMs on GII”—this paper not only explores how to optimize industrial applications using large models but also investigates how these applications can in turn refine the models themselves. This bidirectional interaction perspective significantly enriches the existing literature.
$\bullet$ Cross-Domain Multimodal Integration: Unlike most existing studies that focus solely on specific types of large models, such as language models or vision models, this review encompasses LLMs, vision transformers, multimodal models, and time series models. This cross-modal integration enhances the design, development, and evaluation of complex instrumentation and measurement systems used in generating, acquiring, conditioning, and processing signals. By leveraging the unique advantages across different models, it offers a more comprehensive and in-depth perspective for advancing measurement science and applications, thereby addressing complex industrial challenges more effectively.
$\bullet$ Practical Applications of Continual Learning: This paper emphasizes the real-world application of CL strategies in IIoT-enabled industrial systems, particularly in edge and cloud computing collaborative environments. This focus ensures that the models are not only adaptable to new data and changing conditions but are also resource-efficient. By reducing computational demands and training costs, our approach addresses critical constraints in industrial applications.
II Preliminary
II-A Research Roadmap
Illustrated in Fig. 4, GII has evolved through four crucial stages:
$\bullet$ Traditional Machine Learning, characterized by task-specific algorithm design optimized for small data sets, relying heavily on manual feature extraction and fine-tuning. Representative methods include kNN [37] and SVM [38].
$\bullet$ Neural Networks, marked by enhanced capabilities in handling large and complex data sets, with deep neural networks learning data features automatically, significantly improving pattern recognition accuracy and generalization. Representative models include CNN [39] and ResNet [40].
$\bullet$ Pre-trained Models, where models are pre-trained on massive data sets, enabling fine-tuning on relatively more minor specific tasks to achieve good performance, significantly reducing the time and resources needed for training task-specific models. Representative models include ViT [19] and CLIP [20].
$\bullet$ CLMs, where models can continuously learn in changing tasks and environments, significantly expanding the boundaries of machine learning applications, marking a step towards higher levels of GII. Representative models include GPT [18] and LLaMA [16].
Figure 4: A roadmap for general industrial intelligence with four generations of models based on task solving capabilities.
II-B Industrial Large Models
Industrial LMs are large foundational models pre-trained on diverse industrial data sources, covering various data granularities and modalities, as depicted in Fig. 5.
II-B1 Model Architectures
Transformer [15] is a deep learning model known for its proficiency in processing sequential data types like text or time series data. Its workflow can generally be divided into several steps. Initially, input data is transformed into a set of embeddings, capturing initial data characteristics. Next, the embeddings go through a series of encoders and decoders. At each step, an attention mechanism enables the model to focus on different input data segments. Finally, the model generates output using auto-regressive decoding. Specifically, a brief overview of the architectures in Transformer is provided below.
Embedding. Embeddings convert input data (e.g., words, sentences, or any form of raw data) into fixed-length vector representations. Commonly used tokenizers, such as wordpieces and byte-pair encoding, are frequently employed in this process [41]. These vectors capture essential input features and are subsequently processed by the model. In addition to embedding the input data, the Transformer also integrates positional encoding to maintain the sequential order of words in the sequence. Vanilla Transformer employs Sinusoidal positional encoding [15] or Learned positional encoding [42].
Attention Mechanism. The attention mechanism, which serves as the core of Transformer, enables the model to concentrate on various segments of the input sequence, thereby improving its comprehension of sequence relationships. This process is achieved by calculating weight scores for each input position and using them to generate a weighted average output. The most basic form of the attention mechanism (scaled dot-product attention) can be expressed by
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V,$ (1)
where $Q$, $K$, and $V$ represent the query, key, and value matrices, respectively, and $d_{k}$ denotes the dimension of the key. Transformer utilizes three distinct types of attention mechanisms, each focusing on different inputs: i) Self-Attention: $Q$, $K$, and $V$ all come from the same sample. ii) Encoder-Decoder Attention: The decoder allows focusing on the relevant positions in the input sequence. $Q$ is derived from the output of the prior decoder layer, while $K$ and $V$ come from the encoder’s output. iii) Masked Self-Attention: The decoder’s self-attention layers apply a mask to prevent attending to future positions, which is achieved by setting a part of the input to infinity before the self-attention softmax step.
Figure 5: Industrial LMs are pre-trained on multi-source, multi-granularity and multimodal industrial data and can be adapted for a variety of downstream applications.
Multi-Head Attention. Multi-head attention enables the model to capture information from different representation subspaces simultaneously. In practice, this means passing $Q$, $K$, and $V$ through different linear transformations $h$ times, where $h$ is the number of heads, applying attention mechanisms separately, and subsequently concatenating the outputs. This allows the model to capture different aspects of the sequence at various positions, thereby enhancing the model’s information-processing capabilities. Multi-Head Attention can be expressed by
$\text{MultiHead}(Q,K,V)=\text{Concat}(\text{head}_{1},\dots,\text{head}_{h})W^{O},$ (2)
where each $\text{head}_{i}$ is
$\text{head}_{i}=\text{Attention}(QW^{Q}_{i},KW^{K}_{i},VW^{V}_{i}),$ (3)
where $W^{Q}_{i}$, $W^{K}_{i}$, and $W^{V}_{i}$ are parameter matrices for each head $i$, with $W^{O}$ as the output parameter matrix, and $h$ representing the number of heads. By integrating these elements, the Transformer can process sequential data efficiently and effectively, capturing long-distance dependencies while maintaining relatively low computational complexity. This has made it the architecture of choice for various NLP tasks.
Encoder-Decoder Architecture. In Transformer, the encoder and decoder are separate, each comprising multiple identical layers. Each layer in the encoder contains two sub-layers: a multi-head self-attention mechanism and a simple, position-wise, fully connected feed-forward network. Both sub-layers are equipped with residual connections, followed by layer normalization. The decoder closely resembles the encoder but incorporates two multi-head attention mechanisms instead of one. The first multi-head attention mechanism is masked to avoid positions from attending to the future. An additional sub-layer is inserted between the encoder and decoder to pass the key and value information from the encoder to the decoder.
Auto-regressive Decoding and KV Cache. Auto-regressive decoding generates model outputs sequentially, using previously generated outputs as input for the subsequent step. This method relies on earlier outputs to generate subsequent ones. In Transformer, a key-value cache [43] mechanism can enhance the efficiency of the decoding process by storing and reusing previous attention calculations, thus speeding up generation and increasing efficiency.
We introduce a data-centred taxonomy of industrial LMs. We classify existing LMs into four categories: language-based models, vision-based models, time series-based models, and multimodal models.
II-B2 Large Language Models
LLMs are predominantly constructed using the Transformer [15] architecture and encompass pre-trained LMs whose parameter counts span billions to trillions. These models, extensively pre-trained on vast textual corpora, have amassed a considerable depth of knowledge, empowering them to handle a variety of NLP tasks adeptly. Such tasks notably include text generation, semantic understanding, and sentiment analysis.
LLMs are categorized into open-source and closed-source categories. Open-source models like LLaMA [16], Vicuna [17], and Gemma provide the research community and developers with flexibility by allowing direct access to the model architecture and pre-trained weights. This access facilitates fine-tuning for specialized tasks, showcasing their robust performance in few-shot and zero-shot learning contexts. These contexts highlight their adaptability to new tasks using minimal or no new training data.
Conversely, closed-source LLMs, such as PaLM [44] and GPT-4 [18], are primarily available through API services. These services require users to formulate precise prompts to direct the models’ inference tasks. Although this modality restricts direct control over and transparency of the models’ internal workings, it offers a user-friendly interface advantageous for commercial applications and product integration, enabling effective utilization of these powerful computational tools.
II-B3 Vision-based Models
Early models in computer vision primarily used CNNs like ResNet [40], pre-trained on datasets such as ImageNet, to enhance performance on downstream tasks. The introduction of Vision Transformer (ViT) [19] marked a significant shift by applying the Transformer architecture to image patches, enabling the model to recognize image patterns without the traditional CNN inductive biases like local receptive fields and translational invariance. ViT depends heavily on extensive pre-training to extract discriminative features that capture both local and global image details.
Recent developments have introduced Vision Foundation Models (VFMs) that broadly fall into two categories: language-augmented and purely visual models. Language-augmented VFMs, like CLIP [20], utilize contrastive learning to align text and image data, improving image encoding robustness. Purely visual VFMs, trained with techniques like masked autoencoders, focus solely on visual information, showing strong generalization in various vision tasks and notable zero-shot capabilities.
SAM [45] integrates image and prompt encoders with a mask decoder, all pre-trained on large datasets. It uses textual or spatial prompts to dynamically generate segmentation masks, enabling effective zero-shot adaptation across varying data distributions through prompt engineering.
II-B4 Time Series-based Models
Initially applying LLMs to generic time series forecasting, a one-size-fits-all approach [21] introduces a unified framework using partially frozen LLMs. Specifically, fine-tuning is limited to the embedding and normalization layers, while the self-attention and feedforward layers are frozen. This strategy achieves state-of-the-art or comparable performance across major time series analysis tasks, including time series classification, short/long-term forecasting, interpolation, anomaly detection, and few-shot and zero-shot predictions.
Recent work like Time-LLM [46] proposes reprogramming time series with the source data modality and using natural language prompts to harness the potential of LLMs as effective time series models. It achieves state-of-the-art performance across various prediction scenarios and excels in few-shot and zero-shot settings. Time-LLM is lightweight and efficient as it does not directly manipulate input time series data or fine-tune the underlying LLM. In the industrial field, a method for fault diagnosis using time series data was proposed. This method converts 1-D time series signals into grayscale images using gray-transform and transform-gray coding methods for intuitive fault representation.
Figure 6: (A) Actual bearing fault testbed. (B) indicates four common types of bearing faults1, i.e., (a) Rolling body wear, (b) Inner race wear, (c) Outer race crack, (d) Outer race wear.
11footnotetext: This image is referenced from literature [47].
II-B5 Multimodal Models
Multimodal models typically integrate various types of inputs, such as text, images, and audio, within a single framework for data processing and analysis. By harnessing the fusion of multiple perceptual information, these models aim to enhance task performance accuracy and efficiency.
Multimodal Large Language Models (MLLMs) combine the linguistic processing capabilities of LLMs with functionalities for processing other modalities, making them particularly adept at understanding and generating content involving complex data types, such as text and images. Notable examples include:
$\bullet$ BLIP [48] is designed to enhance image description and comprehension. Utilizing pre-training techniques such as contrastive learning. BLIP-2 [49] employs a dual-stream Transformer structure that optimizes the processing of visual and textual streams. With an improved cross-modal attention mechanism, BLIP-2 achieves more precise alignment between images and text, enhancing the model’s ability to describe and understand image content.
$\bullet$ MiniGPT-4 [50], a scaled-down version of the GPT, focuses on small-scale but efficient image and text processing tasks, maintaining the generative capabilities and flexibility of the GPT series.
$\bullet$ LLaMA-Adapter [51] series and LLaVA [22] utilize adapter modules or simple linear layers to merge visual features with LMs, aimed at minimizing the interference between visual and textual data during training while optimizing the model’s capacity to understand and generate visual information.
TABLE II: The five core subtasks of fault diagnosis
Subtask Description
Detection Determine the presence of a fault.
Classification Identify the type of fault present.
Severity Assess the severity of the fault.
RUL Predict the time until maintenance is required.
Fault tracing Investigate the underlying cause of the fault.
II-C Key Tasks of GII
We categorize industrial tasks into four primary types based on their nature and specific requirements: classification, detection/localization, planning & control, and question & answer (Q&A). Each task type presents unique challenges and demands specialized technological approaches, particularly leveraging IIoT, edge computing, and advanced communication technologies. Detailed discussions on the applications and solutions tailored to each task type will be provided in the subsequent sections.
II-C1 Classification
Classification tasks form the backbone of advanced analytics in industrial systems, especially when integrating IIoT data with LMs for enhanced measurement and decision-making accuracy. These tasks encompass multiple sub-domains, including industrial fault diagnosis and industrial human behavior recognition. LMs, with their robust data processing capabilities and advanced learning algorithms, significantly enhance the accuracy and efficiency of classification tasks.
Mechanical Fault Diagnosis. The fault diagnosis tasks map measures signal features relative to the state of mechanical components [13, 12, 13, 52]. This process involves determining the basic condition of whether a fault exists. It may include more complex estimations such as the fault level, the severity of the fault, the remaining useful life, and the specific cause of the fault. Table II details these five essential sub-tasks of fault diagnosis. Equipment fault diagnosis demands high quality and quantity of data, especially when dealing with different types of equipment and their respective fault types, where relevant data are often scarce. Although there are some open-source datasets available, such as CWRU [53], and XJTU-SY [54], which provide some support for model development, they still fail to fully meet the data requirements of practical applications. Fig. 6 illustrates a typical diagnostic data collection device capable of collecting essential operational data from equipment during actual operations.
Although LLMs such as ChatGPT [18] offer extensive knowledge bases and proficient language processing capabilities, their application in specialized domains like mechanical fault diagnosis is limited due to a lack of domain-specific expertise. Wang et al. [55] propose overcoming these constraints by integrating Local Knowledge Bases (LKB) with LLMs. This integration involves preparing the LKB, vectorizing its content, and applying prompt engineering to tailor the LLM’s responses for domain-specific tasks.
Activity/Gesture Recognition. In automated systems, human activity, state monitoring, human pose estimation and gesture recognition technologies are crucial for quality control, safety monitoring, and process optimization. Typically, data is collected using flexible electronic skin or electromyography (EMG) sensors (see Fig. 7). Flexible electronic skin detects changes in pressure and stress distribution on the skin surface, while EMG sensors monitor the electrical signals generated by muscles during activity.
Figure 7: EMG/E-skin based human activity recognition and gesture recognition.
Flexible electronic skin, placed on key body parts, collects real-time physiological signals processed by deep learning models like CNNs for activity classification. Gesture recognition via surface electromyography involves attaching electrodes to the wrist, capturing, amplifying, filtering, and processing signals through deep learning to translate them into gestures, supporting human-computer interaction. Key components include electrodes, control boards, power units, and communication modules.
II-C2 Detection/Localization
Detection and localization tasks primarily focus on identifying and precisely locating anomalies, defects, or specific features within industrial systems, which are crucial for ensuring product quality and process efficiency. LMs play a pivotal role in these tasks by enhancing the handling and analysis of measurement data. LMs facilitate sophisticated data preprocessing, enabling cleaner and more structured datasets that are essential for accurate detection. Moreover, their advanced capabilities in feature extraction allow for a deeper understanding of complex patterns within the data. Pattern recognition algorithms within LMs efficiently identify and categorize anomalies, while rigorous result validation processes ensure the reliability of the findings.
Industrial Human Action Recognition. In modern industrial settings, Industrial Human Action Recognition (IHAR) is crucial for enabling effective human-machine collaboration, improving quality control, safety monitoring, and optimizing processes. IIoT-connected cameras and depth sensors provide real-time contextual information critical for recognizing human actions in complex industrial environments. The data collected is processed locally at the edge, enabling immediate responses to safety-critical actions.
Recent advances in this domain leverage 3D-CNNs to interpret dynamic scenes and behaviors from video data. Additionally, techniques integrating Human-Object Interaction detection are increasingly applied to enhance the precision of action recognition. Despite these technological advancements, the application of machine learning in IHAR, especially in industrial environments, is still in its nascent stages. Liang et al. [56] have developed approaches that merge foundational models with lightweight architectures to support low-cost dataset creation and enable real-time IHAR recognition.
Figure 8: Industrial anomaly detection.
Industrial Anomaly Detection. Industrial anomaly detection is geared towards identifying behaviours that deviate from normal operational patterns through continuous production process monitoring [14]. Popular datasets such as MVTec-AD [57] and VisA [58] are extensively used for this purpose. MVTec-AD specifically focuses on surface defect detection across various industrial objects made of different materials and shapes. VisA provides a broader spectrum of visual anomalies, enabling extensive testing of models’ generalization capabilities in complex scenarios. Deploying deep learning techniques on these datasets has demonstrated strong potential in effectively addressing unknown types of anomalies (see Fig. 8).
Laser Powder Bed Fusion. Laser Powder Bed Fusion (LPBF) represents an advanced metal additive manufacturing technology, commonly referred to as 3D printing, as shown in Fig. 9. This process utilizes a laser to selectively melt and solidify metal powder layer by layer, allowing for the fabrication of complex metal parts and the manufacturing of end-use components with intricate structures [59].
Figure 9: Additive manufacturing process monitoring.
In LPBF, the deterioration of internal surface quality can compromise the mechanical performance of the final product and may even lead to processing failures. Ma et al. [60] have designed an online monitoring system that adjusts monitoring parameters to address potential defects during the LPBF process, thus facilitating real-time control and quality improvement. The system incorporates a semantic segmentation model for visualizing surface irregularities and a “surface defect area ratio” metric for quantifying defects.
In addition, Era et al. [61] developed an image segmentation framework incorporating SAM [45] and a multi-point prompt generation scheme based on unsupervised clustering. This framework was applied to real-time porosity segmentation during the LPBF process and achieved high Dice Similarity Coefficients without the need for supervised fine-tuning of the model.
Laser Sintered Processing. Due to their superior mechanical, thermal, and dielectric properties, ceramic materials have become crucial in aerospace facilities, biomedical implants, and optoelectronic devices. Traditional ceramic processing methods face numerous challenges, particularly in material densification and dimensional accuracy. Laser sintering has emerged as a powerful tool for enhancing the fabrication of advanced ceramics, offering high resolution, rapid processing cycles, and flexible control [62].
However, the Gaussian distribution of laser energy leads to variations in material density and porosity during the sintering process, potentially reducing the finished products’ mechanical strength and thermal stability. Therefore, accurate evaluation of the material property changes in laser-sintered materials is crucial for optimizing processing parameters to ensure the quality and performance of the end products. As shown in Fig. 10, Wang et al. [63] have developed a micro-pore detection and segmentation model based on Mask R-CNN, providing a precise tool for measuring the surface porosity of laser-sintered alumina ceramics.
II-C3 Planning & Control
The Planning capabilities of LMs are essential for tasks like autonomous driving and robotics, where they optimize decision-making, tool learning [64], route planning, and obstacle avoidance using real-time sensor data. This enables efficient navigation and task execution in dynamic environments. On the Control side, LMs play a key role in communication-control integration, optimizing signal processing, managing network resources, and ensuring reliable data transmission. This is crucial for real-time coordination between edge devices and centralized systems, enhancing both planning and control functions with minimal latency and high accuracy.
Autonomous Driving. Autonomous driving services depends on active sensing from modules such as camera, LiDAR, radar, and communication units [31]. Autonomous driving perception and motion planning integrate generative LLMs with path planning to enable natural language control of autonomous vehicles (see Fig. 11). The process begins with speech-to-text technology converting verbal descriptions into text, which LLMs then interpret to generate executable path planning code. This code is tested in simulation environments, providing feedback for optimizing vehicle operation, enhancing adaptability, and decision-making accuracy [65, 66].
Figure 10: Porosity analysis for femtosecond laser processing.
In dataset developments, the Honda Research Institute’s HAD dataset includes over 5,675 annotated driving video clips [67], while the NuPrompt [68] expands the nuScenes dataset with detailed object-centric linguistic annotations suitable for complex driving scenarios. The NuScenes-QA [69] provides extensive visual scene data for testing visual question-answering capabilities in autonomous driving applications. DriveLM [65], utilizing datasets like nuScenes and CARLA, showcases effective methods for combining graphical visual question answering with end-to-end driving tasks, proving the value of integrating visual reasoning in autonomous vehicle.
Industrial Robotics. In recent advancements, researchers have focused on the integration of LLMs to enhance robotic operations in industrial settings. Fan et al.[70] explore the application of LLMs in the generation, error analysis, and code production of robotic tasks within manufacturing settings.
Industrial Control. Industrial Control Systems control and operate the physical processes of critical national infrastructure sectors, such as water treatment, electricity generation, and manufacturing [71]. Industrial control tasks manage, regulate, and direct systems, equipment, or processes to achieve stable operation and the desired behaviour of the system. Song et al. [72] demonstrated the potential of foundational models in industrial control by employing clustering methods to select exemplars and generate prompts for a base model, which then regulated the air conditioning in buildings accordingly. Joglekar et al. [73] propose a global control strategy based on LLM that allows the transfer of control strategies to a limited set of skills. These skills can execute high-precision tasks through dynamic context switching when specifically trained.
Figure 11: LLM planning and control for autonomous driving.
Co-Design of Communication and Control. As a critical aspect of building the IIoT, the design of communication and control systems has always been a major focus[74]. Due to the strong coupling relationship between the two, it is crucial to develop new metrics that can simultaneously evaluate the performance of both communication and control systems [75]. Wang et al. [76] used the Age of Information (AoI) to characterize “information freshness” and studied the impact of average AoI on control costs. By minimizing AoI, they aimed to reduce the combined costs of control and communication. Chen et al. [77] considered scenarios with multiple users, addressing both uplink and downlink wireless transmissions. Their work integrated multiple antennas on the remote end to improve signal reception quality and developed a loop information age metric. This metric not only captures information freshness but also dynamically allocates the rounds of transmission attempts based on the success or failure of UL and DL transmissions, thereby enhancing communication reliability while ensuring control convergence speed.
As shown in Fig. 12, the co-design of communication and control enhances the stability and responsiveness of IIoT systems by optimizing data transmission and control. LMs, with their strong representation capabilities, offer significant potential in this area. By extracting semantic information [78], LMs can optimize system coordination and improve both communication efficiency and control response, paving the way for more advanced semantic communication and the further development of GII.
II-C4 General Q&A
In the industrial sector, QA systems are crucial for addressing complex queries, supporting decision-making, and improving operational efficiency. By integrating into enterprise knowledge management systems, QA technology helps engineers and operators quickly find solutions, reducing downtime and optimizing maintenance processes.
While LLMs perform well in open-domain QA tasks, their application in specific industrial domains presents challenges. Retrieval-based LLMs often struggle with complex industrial queries requiring deep domain-specific knowledge. Improvement strategies include new interaction paradigms that integrate domain documentation and fine-tuning techniques to provide domain-specific knowledge during runtime (see Fig. 13). Additionally, specialized QA datasets for industrial applications, such as the 32k QA pair dataset in the cloud domain by Wang et al. [79], offer new benchmarks for evaluating and improving LLMs.
Figure 12: Communication and control system.
III Large Models for General Industrial Intelligence
LMs, owing to their extensive internal knowledge and reasoning capabilities, enhance GII in the following two fundamental ways:
(1) As effective enhancers of data, LMs leverage their extensive external knowledge and analytical capabilities to improve the quality of industrial data and optimize existing methodologies;
(2) As efficient predictor in industrial tasks, LMs utilize their rich internal knowledge and emerging inferential abilities to effectively handle various predictive tasks, including both tuning and non-tuning predictors based on whether access to LMs parameters.
III-A LMs as a Data Enhancer
III-A1 Data Denoise
Acquiring large-scale datasets with precise expert annotations in practical settings is costly in terms of time and human resources. Consequently, real-world data, which inherently includes label noise, is commonly used. Addressing this challenge, Cheng et al. [80] propose an adaptive label noise learning framework based on pre-trained models. This framework integrates an adaptive warm-up phase, sample correctness statistics-driven data separation, and a linear decay fusion strategy to handle diverse label noise effectively. Unlike traditional methods targeting specific noise types, this approach enhances learning from noisy data, improving the accuracy and robustness of deep neural networks in real-world applications.
III-A2 Data Selection
Data selection involves identifying subsets of available data that notably enhance model performance. Li et al. [81] propose a quality-based mechanism to evaluate the influence of data on model perplexity, with an emphasis on prioritizing data that demonstrates high utility and training potential. In addition to this, techniques such as data pruning, active learning, and transfer learning are also employed to optimize datasets, all aimed at improving learning efficiency and overall outcomes. In industrial applications, such as robotics [82] and anomaly detection [83], selecting high-quality data is crucial for enhancing operational efficiency and the accuracy of decision-making processes. This adaptive data selection approach supports continuous learning in dynamic environments, allowing industrial systems to maintain both stability and efficiency in the face of production shifts and external influences. Moreover, advanced data selection strategies, such as the importance re-sampling method utilized by Xie et al. [84], further optimize LLMs training by focusing on the most informative samples. This prioritization leads to substantial improvements in training efficiency and contributes to enhanced model performance.
Figure 13: LLMs for industrial question answering.
III-A3 Data Augmentation
The potential of LLMs in data augmentation is prominent, particularly in interpreting industrial time series data through textual descriptions and summaries. SignalGPT, specifically trained for diagnosing electrocardiogram signals, exemplifies this adaptability [85]. Zeng et al. [86] demonstrate the use of Textual Entailment models in zero-shot event detection, enhancing performance through initial large-scale text annotation and subsequent model fine-tuning.
LLMs also excel in integrating diverse industrial data sources, enhancing data processing robustness and decision-making. For instance, Hou et al. [87] leverage GPT-4 to refine hierarchical feature definitions, reducing manual annotation needs and bolstering model stability during testing (see Fig. 14).
III-A4 Data Generation
Deep learning tasks, traditionally supervised, heavily rely on costly and time-consuming labeled data annotation. To address this, current research explores zero-shot learning through generative dataset creation, leveraging widely available data across domains [88, 89].
As shown in Fig. 16, common generative models include Variational Autoencoders (VAEs), Generative Adversarial Networks (GANs) [90], and Diffusion Models. Diffusion models, foundational in generative modeling, demonstrate efficacy across diverse modalities including images, audio, and video.
Specifically, diffusion probabilistic models like Denoising Diffusion Probabilistic Models (DDPMs) [91] and Stochastic Differential Equation (SDE)-based models [92] employ a dual-phase optimization approach. They incorporate a diffusion phase introducing noise and a denoising phase reverting this noise to generate data aligned with the training distribution. DDPMs operate in fixed, discrete steps, iterating noise addition and reversal, while SDE-based models offer continuous flexibility in handling the diffusion process, enabling sample generation at arbitrary points.
Specifically, diffusion probabilistic models like Denoising Diffusion Probabilistic Models (DDPMs) [91] and Stochastic Differential Equation (SDE)-based models [92] employ a dual-phase optimization approach. They incorporate a diffusion phase introducing noise and a denoising phase reverting this noise to generate data aligned with the training distribution. DDPMs operate in fixed, discrete steps, iterating noise addition and reversal, while SDE-based models offer continuous flexibility in handling the diffusion process, enabling sample generation at arbitrary points. In industrial applications where data generation is required on demand, these models prove especially valuable [93]. By generating synthetic data that mirrors real-world conditions, DDPMs and SDE-based models provide the ability to quickly respond to evolving industrial needs, whether in simulation, testing, or adaptive system training.
Furthermore, we have included a comparison between raw and generated data, as illustrated in Fig. 15, highlighting the ability of DDPM to produce samples closely resembling the original time series data.
Figure 14: The generation of unique domain descriptions.
With the increasing demand for AIGC [7], the deployment of AIGC, such as ChatGPT and Dall-E, has garnered widespread attention in mobile edge networks. These networks can provide personalized and customized AIGC services in real time while ensuring user privacy. However, to achieve these capabilities, it is essential to thoroughly explore the background of generative models, the lifecycle of AIGC services, and the cloud-edge-mobile collaborative required to support these services.
Beyond data generation, diffusion models find application in automating dataset construction. Tools like FABRICATOR [94] use them to produce labeled data for NLP tasks, while DYNOSAUR [95] automates training instruction generation based on existing dataset metadata. In visual content generation, diffusion models empower the creation of intricate layouts from textual descriptions, enhancing user control without exhaustive manual inputs.
Figure 15: Comparison of raw time series samples and generative samples using DDPM [91].
III-A5 Reprogramming
Due to various challenges faced by deep learning in resource-constrained domains, such as limited data availability, constrained model development costs, and insufficient pre-trained models for effective fine-tuning, Chen [96] proposes model reprogramming as a solution to bridge this gap. As shown in Fig. 17, model reprogramming enables resource-efficient cross-domain machine learning by repurposing pre-trained models from a source domain to solve tasks in a target domain without fine-tuning, even in potentially significant differences between the domains. In many applications, model reprogramming surpasses transfer learning and training from scratch in performance.
Discussion. LMs, effectively mitigate the sparsity and noise inherent in industrial data as data enhancers, endowing existing general industrial intelligence models with enriched external knowledge and augmented analytical capabilities. This technology’s plug-and-play nature offers flexibility in addressing real-world challenges associated with industrial data and models. Nonetheless, a significant barrier to utilizing LMs as enhancers is the substantial time and cost of processing large-scale datasets. Moreover, the inherent diversity of industrial data and the breadth of application scenarios add complexity to creating universally effective enhancers for various contexts.
LMs as data enhancers represent a promising new direction for augmenting industrial data and models, meriting further exploration in future research. Subsequent studies should concentrate on developing efficient, dependable, and universally applicable plug-and-play solutions that address real-world issues like data sparsity and noise while considering the time and cost efficiency for large-scale dataset applications.
III-B LMs as a Tuning Predictor
During the pre-trained phase, LMs typically acquire the general capability to handle a variety of tasks. However, to effectively deploy these pre-trained models in industrial environments, it is essential to fine-tune them according to specific task requirements, the evolving data distribution, and diverse user preferences. Fine-tuning involves further training the pre-trained model on a specific dataset to better adapt it to the demands of a particular task or domain. This process not only facilitates the infusion of domain-specific knowledge but also aligns the model with specific task instructions.
Figure 16: Different types of generative models.
Several fine-tuning strategies have been developed to meet these needs. Here, we focus on two primary adaptation methods: alignment fine-tuning and efficient fine-tuning. Alignment fine-tuning is designed to adjust the model to better match the semantics and operational standards of specific tasks, while Efficient fine-tuning aims to reduce the computational resources and time required while maintaining model performance. Each method has its own application scenarios and advantages, providing flexible adjustment tools for the practical deployment of LMs in industrial settings.
III-B1 Alignment Fine-tuning
In this section, we introduce Alignment fine-tuning, a fine-tuning approach that enhances the alignment of pre-trained LMs with specific task instructions. Alignment fine-tuning primarily consists of three sub-strategies: Instruction Tuning, Model Refinement, and Model Alignment.
$\bullet$ Instruction Tuning. Instruction Tuning (IT) is a crucial technique for adapting pre-trained LLMs to specific tasks and user preferences in various domains, ensuring alignment with nuanced requirements.
Consider a LM $h(x)$ designed to process input data $x$, which typically comprises commands or queries in natural language. IT is a targeted training approach focused on enhancing the model’s precision and efficiency in handling specific directives. The objective of IT involves the recalibration of $h$ by fine-tuning its parameters utilizing a carefully selected training set $\mathcal{I}={(x_{i},\hat{y}_{i})}_{i=1}^{N}$. In this set, $\hat{y}_{i}$ denotes the anticipated response for the input $x_{i}$, emphasizing improvements in specific tasks or functionalities. The redefined model, noted as $h^{\prime}$, operates under the following rule:
$h^{\prime}(x_{0})=\hat{y}_{0},\quad\forall(x_{0},\hat{y}_{0})\in\mathcal{I}.$ (4)
[…]
The primary distinction between the IT process and general pre-trained lies in the method of loss calculation. Although cross-entropy remains the loss function of choice, during the IT process, the loss is calculated solely based on the machine’s response (i.e., the text following “[BOT]:”). The part corresponding to the user request is excluded from the loss calculation by setting the ignore_index parameter. Fig. 18 shows a training sample of IT.
Figure 17: Model reprogramming.
Faysse et al. [97] provide a detailed exposition of the various evaluation metrics employed during directive fine-tuning for adaptation to downstream tasks, offering practitioners actionable insights for the practical deployment of directive fine-tuned models. By organizing two different types of dialogue task flows, InstrDialog and InstrDialog++, Zhang et al. [98] investigate the problem of Continual Instruction Tuning (CIT). The study demonstrates that sequentially fine-tuning instruction-tuned models yields performance competitive with existing CL methods. Furthermore, rich natural language instructions facilitate knowledge transfer and mitigate forgetting, an advantage that current CL methods scarcely leverage.
$\bullet$ Model Refinement. Like humans, LLMs are prone to errors, such as producing inaccurate translations or retaining outdated information. Ensuring that LLMs remain accurate and up-to-date typically involves direct fine-tuning of the model. While effective, this method can be time-consuming and risk degrading performance on tasks the model has previously learned. To overcome these challenges, the concept of Model Refinement (MR) [99] has been developed as a technique to correct specific errors without disrupting the primary functionalities of the model. This approach, initially guided by a “reliability-locality-efficiency” principle, implements an efficient editing process using gradient descent. Subsequent research [100, 101] have expanded on this foundation, applying gradient decomposition techniques to larger models such as GPT-J-6B and T5-XXL. These methods focus on updating a subset of model parameters to modify the labels of specific inputs and integrate editing through retrieval mechanisms.
Let $h(x)$ be a model that processes input data $x$, such as natural language queries. The model is subjected to an editing set $\mathcal{E}={(x_{e},y_{e},\hat{y}_{e})}_{e=1}^{N}$, wherein $\hat{y}_{e}$ is the correct label for $x_{e}$, yet the model erroneously predicts $y_{e}$. The purpose of MR is to modify $h$ into $h^{\prime}$, improving the model’s accuracy on the editing set $\mathcal{E}$. The goal is for $h^{\prime}$ to accurately predict the correct labels for inputs in $\mathcal{E}$ while preserving its original performance on inputs not included in $\mathcal{E}$. The refined model $h^{\prime}$ functions according to the following definition:
$h^{\prime}(x_{0})=\begin{cases}\hat{y}_{0}&\text{if }(x_{0},\hat{y}_{0})\in\mathcal{E},\\ h(x_{0})&\text{otherwise}.\end{cases}$ (5)
Although these innovations lay a solid foundation for research in Continuous Model Refinement (CMR), the field of continual refinement for LLMs remains largely unexplored. A study has highlighted a potential issue: the optimal location for storing factual data might not coincide with the most effective place for editing it [102]. This discrepancy challenges the traditional ”locate and edit” paradigm and could pose significant concerns in the implementation of CMR. Moreover, questions about whether this approach is suitable for LLMs and whether more memory- and computationally efficient methods can be developed still demand further exploration and resolution.
Figure 18: An example of IT data.
$\bullet$ Model Alignment. Model Alignment (MA) is a fundamental concept in the deployment of AI systems, aimed at ensuring that the actions and outputs of AI models are in harmony with human values, ethics, and preferences. This process involves the use of mathematical models, algorithmic adjustments, and iterative feedback to refine AI behaviours, enabling the model to adapt continuously to a stream of incoming data and remain up-to-date.
Consider $h(x)$ as a model tailored for decision-making scenarios, processing inputs $x$. An alignment dataset $\mathcal{A}={(x_{a},y_{a},\hat{y}_{a})}_{a=1}^{M}$ is utilized, in which $y_{a}$ is the model’s initial decision for the input $x_{a}$, and $\hat{y}_{a}$ is the corrected decision conforming to established ethical standards or desired outcomes. MA focuses on transforming $h$ into $h^{\prime}$, ensuring that $h^{\prime}(x_{a})$ yields $\hat{y}_{a}$ for any $x_{a}$ from the alignment dataset, thus realigning the model’s decisions according to the specified criteria. The adjusted model $h^{\prime}$ operates under the following rule:
$h^{\prime}(x_{0})=\hat{y}_{0},\quad\forall(x_{0},\hat{y}_{0})\in\mathcal{A}.$ (6)
Current research efforts focus on developing methods to minimize the challenges associated with model alignment [103, 104]. These methods implement specific optimization strategies that enhance the adaptability of LLMs, ensuring their outputs remain ethically and factually relevant within dynamic societal and information landscapes.
Furthermore, frameworks for Continuous Model Alignment (CMA), based on Reinforcement Learning (RL) and Supervised Learning (SL), have been proposed. The RL-based CMA framework aims to reduce alignment costs and prevent forgetting by employing more efficient reinforcement learning techniques [105]. The SL-based CMA framework integrates AI systems through direct supervision on dynamic datasets, focusing on maintaining long-term alignment with human values [106]. These frameworks demonstrate how continuous learning and model adjustments can be applied in practical industrial settings, thereby enhancing the operational efficiency and decision-making quality of AI systems.
III-B2 Efficient Fine-tuning
Efficient fine-tuning aims to improve the efficiency of fine-tuning processes for LMs. Illustrated in Fig. 19, efficient fine-tuning methods are categorized as Density-Efficient Fine-Tuning, Memory-Efficient Fine-Tuning, and Parameter-Efficient Fine-Tuning (PEFT).
Figure 19: Illustrations of parameter-efficient fine-Tuning.
Density Efficient. With increasing model sizes, traditional fine-tuning of all parameters becomes increasingly challenging. PEFT methods have gained popularity for effectively adapting pre-trained LMs. Recent research has identified the presence of activation sparsity within the Multilayer Perceptron (MLP) blocks of transformers. Lower activation density enables efficient model inference on hardware optimized for sparsity. Building on this insight, a new method called DEFT [107] has been proposed to reduce loss by promoting higher activation sparsity (or lower activation density) in pre-trained models.
Memory Efficient. Traditional fine-tuning methods rely on backpropagation algorithms. However, as the model size increases, they demand exponentially more memory. Zeroth-order methods theoretically estimate gradients with only two forward passes but are typically very slow when optimizing LMs. MeZO [108] enhances the traditional Zeroth-order Stochastic Gradient Descent method by adapting it to in-place operations, maintaining the memory usage for fine-tuning LMs at the same level as that during inference.
Parameter Efficient. Parameter-efficient fine-tuning methods adapt a large pre-trained LM to specific tasks by fine-tuning only a small subset of parameters while leaving most of the model’s parameters fixed. Below are several common strategies used for parameter-efficient fine-tuning:
  • Adapter Tuning. Adapter Tuning entails adding extra small neural network layers to specific Transformer layers within the pre-trained model. Houlsby et al. [109] design a bottleneck module that initially reduces the dimensionality of the original features, applies a nonlinear transformation, and then restores them to their original dimension. The bottleneck module is integrated into each Transformer layer, usually inserted sequentially after the Transformer layer’s two core components, namely the attention and feed-forward layers, as shown in Fig. 20. During fine-tuning, only the weights of these adapters are updated, leaving the original model parameters unchanged. This effectively directs the model’s learning capacity towards the adapters to facilitate adaptation to new tasks.
  • Prefix Tuning. Prefix Tuning involves adding a series of learnable continuous vectors (referred to as prefixes) to the front of each Transformer layer of the model [110]. These prefixes serve as task-specific signals that guide the model in generating task-oriented responses while keeping most parameters unchanged. Jia et al. [111] introduce Visual Prompt Tuning as an efficient alternative to fully fine-tuning large Transformer models in vision tasks. VPT introduces a small number (less than 1% of the model parameters) of trainable parameters in the input space to adapt to various downstream recognition tasks, achieving significant performance improvements.
  • Prompt Tuning. Prompt tuning primarily adds trainable prompt vectors at the input layer of the model. Specific input texts (i.e., prompts) are chosen or generated using discrete prompt methods to enrich the input text. The augmented input is used to produce model outputs beneficial for downstream tasks [112]. Recent research has also introduced soft prompt methods, which guide frozen LMs in performing specific downstream tasks by directly incorporating prefix prompts into the input [113]. Soft prompts are learned via backpropagation and can be fine-tuned with labelled examples to include signals from end-to-end learning. Wang et al. [114] applied the concept of prompt tuning to CL by creating a prompt pool and fetching prompts relevant to the current task to prevent catastrophic forgetting.
  • LoRA. LoRA [115] accomplishes fine-tuning by altering the weights of the pre-trained model through low-rank decomposition. This technique entails substituting the weight matrices of the model’s linear layers with the product of two low-rank matrices, thereby decreasing the number of parameters requiring updates during fine-tuning. The original weight matrices remain unchanged, with only the introduced low-rank matrices being adjusted during training.
    Figure 20: Adding adapters to Transformers. During training, only newly added Adapters are updated while all the other layers are frozen.
Discussion. This section (including Section III-B1 and Section III-B2) explores the application of LMs as tuning predictors in IIoT-driven general industrial intelligence. By introducing two major strategies—alignment fine-tuning and efficient fine-tuning—it provides an in-depth analysis of methods such as instruction tuning, model refinement, and model alignment, examining how they enhance the adaptability of LMs in various industrial scenarios.
However, challenges in industrial applications remain, particularly in addressing catastrophic forgetting and the high costs associated with training. Future research will need to incorporate advanced optimization strategies, such as in-context learning and chain-of-thought [116] reasoning, to overcome these limitations and ensure stability and reliability in predictions. Moreover, as edge computing and communication technologies continue to advance, the ability to continuously optimize LMs’ performance in dynamic and multimodal data environments will be crucial for the further development of general industrial intelligence.
III-C LMs as a Non-Tuning Predictor
Non-tuning predictors for closed-source models involve structured preprocessing of industrial data to ensure compatibility with the input requirements of LLMs. This process is critical in environments where model parameters cannot be tuned. The process is divided into two main stages:
$\bullet$ Data Preprocessing. Raw industrial data undergoes preprocessing to be suitably formatted for LLMs. This preprocessing typically includes operations such as:
$\begin{split}\hat{X}_{\text{in}}&=\text{Template}(X,P),\\ \hat{X}_{\text{in}}&=\text{Tokenizer}(X),\end{split}$ (7)
where $X$ represents the raw data, and $P$ is the instructional prompt specifying the analysis task’s requirements. The Template(·) structures the data according to predefined formats while the Tokenizer(·) converts the data into a sequence of textual tokens suitable for LLM processing.
$\bullet$ Model Interaction and Response Generation. The preprocessed data, $\hat{X}_{\text{in}}$, is input into the LLM, denoted as $f_{\text{LLM}}$. The model processes these inputs to generate outputs, which are then parsed to extract predictive labels:
$\hat{Y}=\text{Parse}(f_{\text{LLM}}(\hat{X}_{\text{in}})),$ (8)
where $\hat{Y}$ denotes the labels predicted by the LLM based on the processed inputs.
To further enhance understanding, we review prompt-based and tokenizer-based methods that utilize LLMs effectively in non-tuning scenarios.
Figure 21: In-context Learning for classification tasks. Several demonstrations containing both sources (comments for a book) and target classes (sentiments for the book) are fed into the LLM with the test input/query that only includes comment. The LLM can learn from the demonstrations and predicts the result without updating its parameters.
Prompt-based Methods. A notable method translates natural language instructions into executable robot actions [82]. This approach proposes customizable input prompts that seamlessly integrate into robotic execution systems or visual recognition programs, facilitating multi-step task plans while mitigating token limitation constraints of ChatGPT. In autonomous driving, LimSim++ [117] provides diverse types and modalities of prompt inputs, extracting road network and vehicle surroundings information at each decision frame, which is then conveyed in natural language to the driving agent (e.g., GPT-4).
Tokenizer-based Methods. The introduction of LiDAR-LLM [118] leverages LLMs’ reasoning capabilities to comprehend outdoor 3D scenes fully. Addressing the scarcity of paired 3D LiDAR and text data, a three-stage training strategy generates datasets that align the 3D modality with the language embedding space of LLMs. Initially, LLM tokenizers were not designed to handle numerical data or recognize associations in time series effectively. This challenge was addressed through lightweight embedding layers and refined prompt engineering [119]. LLMTime [120] utilizes innovative tokenization techniques, transforming tokens into flexible, continuous-valued representations, enabling unadjusted LLMs to achieve or surpass traditional performance in zero-shot prediction tasks.
Transitioning to in-context learning, we explore its relevance and applications in non-tuning LLM scenarios.
In-Context Learning (ICL). Introduced in GPT-3 [121], ICL, initially popularized by GPT-3, enables LLMs to make accurate predictions using minimal demonstrations, bypassing the need for gradient updates. This update-free feature not only lowers adaptation costs but also broadens the model’s applicability across various tasks. The efficacy of ICL hinges on both the pre-training of the model on extensive datasets and the careful selection of demonstrations for task-specific predictions. Studies have identified several critical factors influencing the performance of ICL, including the label space, input text distribution, and the format of demonstrations [122]. Moreover, structural similarity, diversity, and complexity within demonstrations are also significant [123]. Recent research by Coda-Forno et al. [124] has shown that the capabilities of LLMs in ICL can be recursively enhanced through the ICL process itself. Fig. 21 shows an example of in-context learning.
Discussion. This section explores the application of closed-source large models in industrial environments through structured data preprocessing and model interaction in non-tuning predictor scenarios. As edge computing and communication technologies evolve, effectively leveraging non-tuning capabilities of large models to handle multimodal data and real-time tasks in industrial settings will be critical for optimizing general industrial intelligence systems. Future research will focus on refining prompt engineering and improving tokenization techniques to enhance the adaptability and stability of non-tuning large models in complex industrial scenarios.
IV Large Models on General Industrial Intelligence
In the domain of LMs, traditional pre-trained often relies on processing massive datasets, such as corpora comprising billions of words. Faced with new data, these models typically require retraining from scratch. However, distribution shifts introduced by new data often degrade performance on previous tasks or impede adaptation to new data. Employing CL strategies for large models can significantly reduce computational resource consumption and enhance efficiency.
The essence of CL strategies is to enable large models to adapt to data evolving over time and across domains, thereby maintaining the accuracy and freshness of their knowledge. This strategy provides a robust framework for large models, enhancing their adaptability to real-time changes in industrial environments. Challenges in CL include:
$\bullet$ Catastrophic Forgetting: Models frequently forget previously learned tasks when adapting to new ones, particularly evident in horizontal CL as models switch between tasks with varying data distributions.
$\bullet$ Extended Task Sequences: Facing time-series data or evolving task structures, models require continuous updates, which may lead to forgetting earlier learned tasks.
$\bullet$ Distributional Shifts: Sudden changes in data distribution pose greater demands on the model’s generalization capabilities and adaptability.
$\bullet$ Unknown Task Learning: Incorporating new and previously unseen tasks requires robust mechanisms to adapt without forgetting existing tasks.
$\bullet$ Incorporation of New Modalities: As industrial environments integrate new data modalities (e.g., visual, audio), models need the ability to align and learn from these modalities without disrupting existing knowledge.
Recent research have shown that with simple and scalable CL strategies, large models can be effectively updated using only a fraction of the computational resources, achieving performance comparable to that of full retraining [125]. While these advancements are promising, further exploration is needed to assess their effectiveness, particularly in the context of LLMs applied to industrial environments. This section reviews classic CL methods, investigates their application in pre-trained models, and explores the latest online adaptation techniques relevant to general industrial intelligence.
IV-A Classical Continual Learning
CL aims to enable models to adapt to new information continuously while retaining mastery over previously acquired knowledge, as illustrated in Fig. 22. The objectives of CL include reducing the forgetting of previous knowledge and leveraging prior knowledge to expedite new task learning, thereby improving efficiency and reducing computational resource consumption.
CL is currently used in various industrial domains, such as fault diagnosis [126] and target detection [127]. Empirical studies such as those by Luo et al. [128] highlight the challenges of catastrophic forgetting in LLMs and suggest methods to mitigate these effects through innovative CL strategies. These include regularization methods that aim to preserve previous knowledge by adding constraints during the learning of new tasks, and replay methods that involve reusing old data to reinforce memory. Moreover, parameter isolation techniques provide another avenue for CL by dedicating specific parameters to particular tasks, thus protecting learned knowledge from being overwritten.
In summary, balancing learning plasticity and memory stability is crucial in CL, requiring strategies that can adapt models to new data without losing the ability to perform well on previously learned tasks. As shown in Fig. 23, we mainly discuss classical CL as (1) Regularization-based methods, (2) Replay-based methods, and (3) Parameter Isolation methods.
Figure 22: General framework of CL. New tasks arrive sequentially, and the model needs to learn them incrementally.
Continual LearningReplay-basedRegularization-basedParameter IsolationParameter RegularizationFunctional RegularizationExperience ReplayGenerative ReplayStatic ArchitectureDynamic Architecture EWC [129], Online EWC [130], SI [131], MAS [132], TFCL ([133, 134, 135]), SPLR [136], KCL [137], ACML [138], RTRA [139], CoSR [140] LwF[141], ANCL [142], TAM-CL [143], PRD [144], TARGET [145] iCaRL [146], TinyER [147], DER [148], EPR [149], AdaER [150], GEM [151], A-GEM [152], OCD-Net [153], CRL [154], PODNet [155], DER/DER++ [148], D3former [156], AAER [157], EPR [158], CPL [159], RR [160], InfluenceCL [161], BFP [162], LP-DiF [163] DGR [164], DGM [165], FearNet [166], SRM [167], EEC [168], MAE [169], DDGR [170], CPDM [171], DiffusePast [172], FSIL-GAN [173] DSG [174] PathNet [175], PackNet [176], HAT [177], RPS-Net [178], UCL [179], CLNP [180], CDG [181], DSN [182] PNNs [183], Expert Gate [184], CCLL [185], PI-GNN [186], DEGC [187], SEDEM [188], SANE [189], Dynamos [190], Expo [191]
Figure 23: Taxonomy of CL.
IV-A1 Regularization-based Methods
Regularization-based methods commonly employ regularization terms to prevent catastrophic forgetting during training. According to the regularization target, these methods can be further categorized into two sub-directions: parameter regularization and functional regularization.
Parameter Regularization Methods. In parameter regularization methods, regularization terms constrain the variations of parameters to balance the plasticity-stability trade-off. The regularization term is calculated to discourage significant changes in these critical parameters while training new tasks, as shown in Fig. 24. For task $t$ with the parameter $\theta$, the loss function ~$\tilde{\mathcal{L}}_{t}(\theta)$ is [131]
~$\tilde{\mathcal{L}}_{t}(\theta)=\mathcal{L}_{t}(\theta)+c\sum_{i}\Omega_{i}^{t}\left(\theta_{i}-\theta_{t-1,i}^{*}\right)^{2},$ (9)
where $\mathcal{L}_{t}(\theta)$ is the task-specific loss for task $t$, $i$ labels each parameter, $\Omega^{t}_{i}$ is the importance matrix of the $i$-th parameter when learning task $t$, $\theta_{t-1,i}^{*}$ is the $i$-th parameter after learning task $t-1$, and $c$ is a hyperparameter that determines the importance of the previous task compared to the new one. As the first work proposing the regularization-based method, Elastic Weight Consolidation (EWC) [129] estimates the importance of each parameter based on its significance for the previous tasks. Consider the training from a Bayesian framework [192], the conditional probability $p\left({\theta}\mid\mathcal{D}_{1:t}\right)$ after learning the task $t$ can be obtained by [193]
$\centering\begin{split}p\left({\theta}\mid\mathcal{D}_{1:t}\right)&\propto p({\theta})\prod_{i=1}^{t}p\left(\mathcal{D}_{i}\mid{\theta}\right)\\ &\propto p\left({\theta}\mid\mathcal{D}_{1:t-1}\right)p\left(\mathcal{D}_{t}\mid{\theta}\right),\end{split}\@add@centering$ (10)
where the distribution $p\left({\theta}\mid\mathcal{D}_{1:t-1}\right)$ incorporates all information from previous tasks [129]. Typically, the posterior distribution $p\left({\theta}\mid\mathcal{D}_{1:t-1}\right)$ is intractable, necessitating approximation (e.g., Laplace approximation, Variational inference, and Moment matching). In Laplace approximation [192], the posterior distribution is treated as a multivariate Gaussian, i.e., $p(\theta\mid\mathcal{D}_{1:t-1})\approx q_{t-1}(\theta)\approx\mathcal{N}\left(\theta;\theta_{t-1}^{*},\Lambda_{t-1}^{-1}\right)$, where $\Lambda_{t-1}$ is the precision matrix. To calculate the precision matrix, EWC applies Fisher information matrix (FIM) with a diagonal approximation to approximate the Hessian matrix. Then, the loss function when learning task $t$ in EWC becomes
~$\centering\tilde{\mathcal{L}}_{t}(\theta)=\mathcal{L}_{t}(\theta)+c\sum_{i}F_{i}\left(\theta_{i}-\theta_{t-1,i}^{*}\right)^{2},\@add@centering$ (11)
where $F_{i}$ is the FIM of the $i$-th parameter.
Besides Laplace approximation, Variational inference (VI) [193, 194] is also commonly employed for approximating the posterior distribution. VI utilizes optimization techniques to estimate the intractable probability distributions. It aims to determine the distribution that best approximates the actual one, typically by using the KL divergence as the optimization criterion [195]:
$\centering q_{t}({\theta})=\arg\min_{q\in\mathcal{Q}}\operatorname{KL}\left(q({\theta})\|\frac{1}{Z_{t}}q_{t-1}({\theta})p\left(\mathcal{D}_{t}\mid{\theta}\right)\right),\@add@centering$ (12)
where $\mathcal{Q}$ is the set of allowed approximate posteriors, $q_{0}(\theta)=p(\theta)$, and $Z_{t}$ is a normalizing constant. Variational Continual Learning (VCL) [195] combines online variational inference (VI) with Monte Carlo VI [196] to construct a general framework for both deep discriminative models and deep generative models. The VI formulation naturally includes a regularization term. In VCL, training in Eq. 12 is equivalent to maximizing the variational lower bound to the negative online variational free energy or the online marginal likelihood, which is
$\centering\mathcal{L}_{t}\left(q_{t}({\theta})\right)=\mathbb{E}_{{\theta}\sim q_{t}({\theta})}(\ell_{t}(\theta))-\mathrm{KL}\left(q_{t}({\theta})\|q_{t-1}({\theta})\right),\@add@centering$ (13)
where the KL-divergence can be computed in closed form.
Figure 24: Illustration of the effect of parameter regularization across two tasks: $\theta_{1}$ represents the parameter after learning task 1. $\theta_{2}$, $\dot{\theta}_{2}$, and $\ddot{\theta}_{2}$ represent the parameters after learning task 2 under three conditions: parameter regularization, $L_{2}$ regularization, and no regularization, respectively. While $\theta_{2}$ learns task 2 while still retaining knowledge of task 1, $\dot{\theta}_{2}$ learns only task 1 under the overly stringent constraint, and $\ddot{\theta}_{2}$ learns only task 2, forgetting task 1 without any penalty [129].
Considerable efforts have been committed to enhancing parameter regularization approaches. SI [131] performs the importance measurement by considering the parameter’s contribution to the loss decay in an online manner and along the entire learning trajectory. MAS [132] accumulates the importance measure for each parameter via the sensitivity of the predicted output function to changes based on unlabeled data in an online manner. TFCL [133] combine MAS with online CL to extend task-based sequential CL into task-free settings, where data distributions change gradually and the concept of distinct tasks is absent. SPLR [136] learns to task invariant representation and task-specific features across the entire parameter trajectory via SI.
For those methods that employ VI strategies, building upon VCL, CLAW [197] adapts the local and global components of the model architecture with an attention mechanism. For each neuron, one binary parameter controls whether adaptation occurs, along with two parameters regulating the magnitude of adaptation. KCL [137] utilizes an episodic memory unit to store a subset of samples for each task, enabling the learning of task-specific classifiers through kernel ridge regression. ACML [138] establishes a one-to-many mapping between a task and meta-knowledge components to facilitate meta-knowledge sharing among heterogeneous tasks through an Indian Buffet Process prior [198].
Functional Regularization Methods. Most functional regularization methods are based on knowledge distillation. Their key strategy to prevent catastrophic forgetting is recording the network outputs of a new task before its training, then utilizing these recorded outputs to distil the previously learned knowledge [199]. One classical functional regularization method is LwF [141]. The learning procedure for each task in LwF is divided into two stages: initialization and training. During the initialization stage, the output of the previous networks for the new task is computed and recorded, and the task-specific parameters for a new task are randomly initialized. During the training stage, two losses are utilized to update parameters: one for new specific tasks ($\mathcal{L}_{\text{n}}$) and another for distillation ($\mathcal{L}_{\text{o}}$). For new tasks, the loss $\mathcal{L}_{\text{n}}$ is calculated based on the ground truth $\mathbf{y}_{n}$ and the predictions $\hat{\mathbf{y}}_{n}$. For each original task, the knowledge distillation loss $\mathcal{L}_{\text{o}}$ is applied to ensure that the output probabilities closely align with the recorded outputs from the original network. As for multi-class classification with $l$ labels, these two loss can be estimated by
$\centering\mathcal{L}_{\text{n}}\left(\mathbf{y}_{n},\hat{\mathbf{y}}_{n}\right)=-\mathbf{y}_{n}\log\hat{\mathbf{y}}_{n},\@add@centering$ (14)
$\centering\mathcal{L}_{\text{o}}\left(\mathbf{y}_{o},\hat{\mathbf{y}}_{o}\right)=-\sum_{i=1}^{l}y_{o}^{\prime(i)}\log\hat{y}_{o}^{\prime(i)},\@add@centering$ (15)
where $y_{o}^{\prime(i)}$ and $\hat{y}_{o}^{\prime(i)}$ are modified by softmax temperature ($T=2$ in LwF) [200] from the recorded probabilities $y_{o}^{(i)}$ and current probabilities $\hat{y}_{o}^{(i)}$. Thus, the optimization objective for the shared parameters $\theta_{s}$, task-specific parameters for previous tasks $\theta_{o}$, and the new task-specific parameters $\theta_{n}$ is as follows:
[…]
where $\lambda$ is a hyperparameter that determines the weight of the previous task performance and $\mathcal{R}$ is a regularizer.
A growing number of functional regularization variants have been developed. EBLL [201] constructs a task-specific under complete autoencoder for each task to capture the critical features from previous tasks. It then ensures that the reconstruction of these features from changing is prevented when learning a new task. Based on LwF, LwM [202] proposes an attention-based method to transfer knowledge and restrict the divergence between student and teacher in CIL problems. DMC [203] introduces a double distillation loss to mitigate the bias caused by information asymmetry between previous and new classes. AYT [204] introduces Teacher Adaptation, allowing teacher and student models to update simultaneously when learning new tasks in exemplar-free CIL settings.
IV-A2 Replay-based Methods
Replay-based methods prevent catastrophic forgetting by replaying a part of data from previous tasks and using them when learning a new task. Depending on the way to replay the data, replay-based methods can be further categorized into two classes: experience replay/rehearsal and generative replay/pseudo-rehearsal.
Experience Replay Methods. In experience replay methods, the model trains using a part of previous data samples stored in a memory buffer B and the new task data to prevent catastrophic forgetting (see Fig. 25). Based on LwF, iCaRL [146] proposes this idea to tackle CIL problems. It sets the exemplar sets $\mathcal{P}$ that updates whenever learning new classes. Here, a prioritized list $P$ represents the exemplar set from a class, and $p$ represents a specific exemplar within $P$. $P$ is defined as $P=\left(p_{1},p_{2},...,p_{m}\right)$. The examples at the beginning of the set are more important than those that follow. $m$ denotes the capacity of an exemplar set, i.e., the number of exemplars it can store. If $K$ denotes the total number of exemplars that can be stored and $t$ is the number of observed classes, then $m$ can be calculated as $m=K/t$. For $i=1,2,...,m$ in the exemplar set selection, the exemplar $p_{i}$ is selected by
$\centering p_{i}\leftarrow\underset{x\in X}{\operatorname{argmin}}\left\|\mu-\frac{1}{i}\left[\varphi(x)+\sum_{j=1}^{i-1}\varphi\left(p_{j}\right)\right]\right\|,\@add@centering$ (17)
where $X=\left\{x_{1},x_{2},...,x_{n}\right\}$ denotes the data set of a class, $\varphi(\cdot)$ is the feature function, and $\mu$ is the class mean that calculated by $\mu=\frac{1}{n}\sum_{x\in X}\varphi(x)$. Upon starting to learn class $t$, iCaRL possesses the exemplar set $\mathcal{P}=\left(P_{1},P_{2},...,P_{t-1}\right)$, containing the exemplars selected from class $1$ to class $t-1$.
In addition to using herding selection as iCaRL does [205, 155], various methods can be used, e.g., reservoir sampling [147, 148, 153, 206], gradient-based sampling [207], ring buffer [151], k-Means [147], etc. Moreover, various methods for enhancing sampling strategies are proposed [208, 209, 149, 150]. EPR [149] selects the previous data through saliency maps that provide visual explanations for the model predictions. AdaER [150] selects examples based on contextual cues that consider the interference caused by the data-conflicting and the task performance.
Instead of improving sampling strategies, some methods prioritize learning more knowledge from the memory buffer. GEM [151] treats the loss for each previous task as inequality constraints to prevent their increase, and violations are addressed using gradient projection. Some advancements apply experience replay combined with other machine learning methods, such as meta-learning [210] and contrastive learning [211, 212, 153, 154, 206, 213]. KD is also widely employed in experience replay [155, 148, 156, 157, 158, 159, 160, 161, 163]. PODNet [155] utilizes spatial-based distillation loss all over the network to constrain representation evolution. AAER [157] utilizes anchors to tackle drift prevention and compensation. BFP [162] introduces a new distillation loss based on feature projection to preserve the linear separability of previous classes while maintaining the plasticity.
Figure 25: Framework of two replay-based methods for learning task $t$. In the experience replay method, the new continual model $M_{t}$ is trained on both the current input $(x_{t},y_{t})$ from task $t$ and the replayed data $(x_{t}^{\prime},y_{t}^{\prime})$. In the generative replay method, the generated input $x_{t}^{\prime}$ is input into the previous continual model $M_{t-1}$. The output $y_{t}^{\prime}$ from $M_{t-1}$, paired with $x_{t}^{\prime}$, is utilized to train the new continual model $M_{t}$ with $(x_{t},y_{t})$. The new generator $G_{t}$ is trained on $x_{t}$ and $x_{t}^{\prime}$.
Generative replay Methods. Generative replay methods train an additional generative model to generate pseudo-data for previous tasks, while experience replay methods require substantial memory resources to store and replay previous data, rendering them potentially impractical in specific real-world scenarios [164, 174]. DGR [164] utilizes GANs as a generator to replay previous data and represents the model $H$ as $\left\langle G,S\right\rangle$, where $G$ is the generator and $S$ is a task solving model. The scholar’s training involves training the generator $G$ and training the solver $S$. The generator is trained on the current input $\bm{x}$ and the replayed input $\bm{x}^{\prime}$. The solver is trained on the current data pair $\left(\bm{x},y\right)$ and the replayed data pairs $\left(\bm{x}^{\prime},y^{\prime}\right)$, where $y^{\prime}$ is obtained by inputting $\bm{x}^{\prime}$ into the old solver. When learning the task $t$, the scholar $H_{t}$ learns both the task $t$ and the knowledge of $H_{t-1}$. The train loss is computed by
$\mathcal{L}_{t}(\theta)$ $=r\mathbb{E}_{(\bm{x},\bm{y})\sim D_{t}}[L(S_{t}(\bm{x}),\bm{y})]$ (18)
$\quad+(1-r)\mathbb{E}_{\bm{x}^{\prime}\sim G_{t-1}}[L(S_{t}(\bm{x}^{\prime}),S_{t-1}(\bm{x}^{\prime}))],$
where $r$ is the mixing ratio of real and replayed samples.
Many methods use GANs as generators [214, 215]. DGM [165] introduces conditional generative networks that learn sparse attention masking for connection weights or layer activations during adversarial training. GFR [216] integrates GANs with a feature extractor for feature replay. GNR [217] utilizes the generated data as negative examples to improve the learning of current class features. GarDA [218] and MFA [219] investigate the utilization of GANs in unsupervised domain adaptation [220].
Autoencoders and VAEs [221] are also utilized for data generation. Earlier approaches include but are not limited to FearNet [166], SRM [167], and EEC [168]. MAE [169] employs masked autoencoders that randomly mask patches of input images for enhancing replay efficiency and stability. More recently, diffusion probabilistic models [222] have been employed as generators [170, 171]. KD is applied to generative replay as well [218, 219, 223, 224].
IV-A3 Parameter Isolation Methods
Parameter Isolation Methods learn specific parameters within a network for each task to prevent catastrophic forgetting. Depending on whether the network structure remains fixed during training, parameter isolation methods can be categorized into static and dynamic architecture.
Static Architecture Methods. Static architecture methods allocate their fixed parameters in limited network architecture to tasks for task-specific learning. PathNet [175] introduces pathways containing active modules in each layer through a network. Each task is assigned its optimal pathway. Once task $t$ has been trained, its optimal fit pathway becomes fixed, preventing further parameter changes. PackNet [176] utilizes pruning to free up parameters for new tasks. During the training of PackNet, a network is learned for the initial task. Next, a portion of the learned weights is set to zero for pruning, and these pruned weights are then learned for the second task. In subsequent training, after the network has learned task $t$, pruning is applied to the weights acquired for that task. The pruned weights then undergo retraining for task $t+1$, while the weights related to tasks prior to task $t+1$ remain frozen to retain previously acquired knowledge. Masks are stored to indicate the relevant active parameters when the network learns a new task. HAT [177] introduces a task-based hard attention mechanism with unit-based masks to retain the previous knowledge and constrain updates for news tasks. More recently, DSN [182] proposes a neuron-wise mask mechanism to select the affiliated weights of a small set of neurons for previous knowledge transfer.
Dynamic Architecture Methods. Dynamic architecture methods expand the network as new tasks are learned while retaining the parameters for previous tasks. PNNs [183] adapt to each task by creating a new network for each one. Lateral connections to features of previous networks are used to transfer previous knowledge. DEN [225] utilizes group-sparsity regularization to determine the addition of neurons at each layer for each task, thereby avoiding unnecessary network retraining. RCL [226] trains a recurrent neural network to determine the optimal architectural hyper-parameters for each task. The recurrent neural network is guided by a reward signal derived from reinforcement learning. Based on Mixture-of-Expert [227], some methods [184, 191] utilize task-specific experts for network expansion. In Expert Gate [184], a new expert model is integrated each time a new task needs to be learned, with knowledge transferred from prior models. Dynamic architecture approaches are also employed in continual graph learning [186, 187].
IV-B Pre-trained Models Based CL
In this section, we will explore Pre-trained Models (PTMs) based CL. Inspired by [228], we further summarize and introduce developments in PTM-based continual learning.
Figure 26: Mixture of Experts.
IV-B1 Mixture of Experts-based Methods
As shown in Fig. 26, the Mixture of Experts (MoE) employs a modular approach to enhance both performance and efficiency in neural networks. This approach consists of multiple expert sub-networks, each specializing in different tasks, coordinated by a routing network. The routing network directs incoming data to the most relevant experts based on data characteristics. These experts process the data independently, and their outputs are often combined through mechanisms like weighted averaging to generate the final response.
The modular structure of MoE reduces the overall parameter load by activating only the necessary experts for each input, making it well-suited for large-scale, diverse datasets. In CL settings, MoE helps mitigate catastrophic forgetting by dynamically adjusting which experts are involved based on the task at hand [229].
Recent CL advancements have integrated MoE with foundational models to boost adaptability and minimize forgetting. For example, MoE adapters in pre-trained models such as CLIP [20] enable incremental learning without the high computational costs usually associated with extensive model updates. These adapters dynamically adjust to new tasks, ensuring sustained performance across evolving learning scenarios.
Innovations like the expert aggregation loss mechanism further enhance model adaptability by fine-tuning expert contributions in real-time, based on confidence levels and specific knowledge requirements, providing an efficient framework for managing diverse data streams [230, 231].
IV-B2 Analytic Learning-based Methods
Figure 27: Analytic CL Based on Pre-trained Models vs. Traditional Replay-based CL. Analytic CL, which is example-free, demonstrates performance comparable to that of replay-based methods.
Training deep models using backpropagation has inherent limitations, such as gradient vanishing/exploding, divergence during iterations, and long training cycles. Analytic Learning (AL) addresses these issues by employing recursive least squares (see Fig.  27). Utilizing the matrix inversion principle, its effectiveness has been preliminarily validated. However, since the weights are learned by processing the entire dataset, AL may encounter memory insufficiency problems. This issue can be mitigated by Block-wise Recursive Moore-Penrose Learning (BRMP) [232], which replaces joint learning with a recursive approach while maintaining equivalent model accuracy.
Analytic Class Incremental Learning (ACIL) [233] is the first method to apply Analytic Learning to Class Incremental Learning (CIL) problems. This method is example-free, providing an effective strategy to counter catastrophic forgetting and reducing resource consumption while achieving performance comparable to replay-based methods. Gaussian Kernel Embedded Analytic Learning [234] further addresses catastrophic forgetting in few-shot scenarios by employing kernel methods.
IV-B3 Prompt-based Methods
VPT [111] applies a prompt-based tuning approach where a set of trainable prompts $P=\{P_{1},P_{2},\ldots,P_{M}\}$ is introduced at the model’s input frontend. Combined with input data, these prompts directly influence the forward propagation process to encode task-relevant information.
The selection and application of prompts are realized through an optimization problem, formulated as:
$\min_{P,W}\sum_{(x,y)\in D}\ell(W^{T}\phi(x;P),y),$ (19)
where $\phi(x;P)$ represents the prompted input features, and $\ell$ is the loss function measuring the deviation between predictions and actual values.
Prompt-based CL methods provide an efficient way to adjust pre-trained models, enabling them to adapt quickly to new tasks while retaining previously learned knowledge. A prompt pool is designed to manage and apply these prompts [114], allowing the model to choose the most suitable prompts for specific instances. The selection process is governed by a retrieval function $g(x)$, which evaluates each prompt’s relevance to the current input $x$. The selection strategy can be expressed as:
$K_{x}=\operatorname{argmin}_{s_{i}\in[1,M],i=1}^{N}\gamma(\phi(x,k_{s_{i}})),$ (20)
where $k_{s_{i}}$ are keys associated with the prompts $P_{s_{i}}$, and $\gamma$ represents the cosine distance, used to assess the similarity between input features and keys.
IV-B4 Representation-based Methods
Pre-trained models are widely used for classifying new tasks due to their strong representational abilities. However, evaluating their inherent classification capacity poses challenges. SimpleCIL [235] offers a straightforward and effective approach by calculating the center or prototype of each class in the data stream:
$c_{i}=\frac{1}{K}\sum_{j=1}^{|D^{b}|}\mathbb{I}(y_{j}=i)\phi(x_{j}),$ (21)
where $\mathbb{I}(\cdot)$ is the indicator function which outputs 1 if the expression holds and 0 otherwise, $K$ is the number of samples for class $i$, and $\phi(x_{j})$ is the embedding of input $x_{j}$. This method extracts each class’s most common representational pattern, providing a strong starting point for subsequent classification tasks.
ADAM [236] enhances adaptability by linking the features of pre-trained models with those of task-specific fine-tuned models. This approach bridges domain gaps, improving the generalizability and task adaptability of representations through parameter-efficient modules:
$c_{i}=\frac{1}{K}\sum_{j=1}^{|D^{b}|}\mathbb{I}(y_{j}=i)[\phi(x_{j}),\phi(x_{j};\text{PEFT})],$ (22)
where $\phi(x_{j};\text{PEFT})$ represents the task-specific fine-tuned model embeddings.
To further enhance classification performance, RanPAC [237] introduces feature projection into a high-dimensional space to increase class separability:
$c_{i}=\frac{1}{K}\sum_{j=1}^{|D^{b}|}\mathbb{I}(y_{j}=i)\text{Proj}(\phi(x_{j};\text{PEFT})).$ (23)
The methods discussed in this section demonstrate how to effectively utilize the inherent representational capabilities of pre-trained models to adapt to new tasks and environments, which is crucial for deploying CL systems in dynamic settings.
IV-C Efficient Inference Algorithm
Figure 28: Test Time Adaptation.
IV-C1 Test Time Adaptation
As shown in Fig. 28, Test-time Adaptation (TTA) adjusts a pre-trained model to unlabeled data from a target domain at the time of testing, differentiating it from Domain Generalization which only operates during training. TTA enhances model performance by utilizing available test data, unlike Domain Adaptation that requires access to both source and target domain data. This makes TTA especially suitable for contexts where data privacy is paramount. TTA methodologies can be divided into test-time domain adaptation, test-time batch adaptation, and online test-time adaptation, based on test data characteristics.
The Edge-Cloud Learning Model [238] adapts to dynamic edge environments through a modular architecture. This approach enables the derivation of personalized edge models from large cloud-based models, tailored to specific needs while minimizing parameter conflicts. This modular setup allows flexible and efficient model adaptation in resource-constrained environments.
Continual Test-Time Adaptation addresses the challenge of adapting models to continuously evolving target domains. The introduction of the Visual Domain Adapter (ViDA) [239] allows the model to handle both domain-specific and shared knowledge. The Homeostatic Knowledge Allotment strategy optimizes the integration and utilization of knowledge from different ViDA, enhancing the model’s adaptability and preserving its plasticity across varying domains.
IV-C2 Edge Cloud Collaboration
Resource constraints in edge environments often hinder local deployment of LMs, leading to significant inference delays [240] that challenge real-time IIoT intelligent perception applications such as autonomous driving, metaverse [241], and digital twins [10]. Edge-cloud collaborative IIoT perception methods, leveraging cooperative inference between large and small models, provide a viable solution to this issue [242, 243]. By employing a hard example mining algorithm, difficult samples are selectively sent to the cloud for processing by a large visual model, while simpler samples are handled locally by small models, thereby enhancing performance on challenging samples while ensuring low inference latency.
As illustrated in Fig. 29, in edge-cloud collaborative intelligent driving systems, cloud-based large models play a critical role, particularly in navigating complex environments with high object density or when local model performance is compromised by visual impairments or lighting variations. Leveraging cloud resources for deeper analysis enhances accuracy, reliability, and safety in these systems.
The Cloud-Edge Collaborative Computing (CECC) framework [244] integrates computing resources from network back-end and edge nodes to deliver more flexible services, striking an optimal balance between resource abundance and responsiveness. Strategies aimed at mitigating edge device limitations emphasize enhancing edge-cloud collaboration. For example, the EdgeFM system [245] augments edge capabilities using cloud servers for querying cloud-based Feature Maps, ensuring real-time model adaptation and maintaining high accuracy under dynamic operational conditions. The EVOLVE framework [230] further advances cloud-edge synergy by employing multiple cloud-deployed expert models to refine self-supervised learning on edge devices. Featuring a dynamic expert aggregation loss mechanism, EVOLVE adjusts model weights in real-time based on reliability and expertise, effectively managing dynamic data streams and sustaining high accuracy in dynamic environments.
Figure 29: Edge-Cloud Collaborative Driving Planning with LMs.
IV-C3 Federated Learning
The rapid development of federated learning (FL) has significantly advanced the progress of GII [246, 247, 248, 249]. FL enables the training of a global model through the exchange of model parameters or gradients between a central server and edge devices, while simultaneously preserving data privacy and effectively mitigating the challenges posed by data silos. This decentralized approach has proven to be a powerful solution in scenarios where sensitive data must remain localized, offering significant advantages for real-world applications.
Research further suggests that the integration of FL with GII holds considerable promise for future advancements [250]. The combination of FL with the IIoT expands application possibilities across multiple domains, such as data sharing, task offloading, attack detection, localization, and privacy/security enhancement [33, 251].
In particular, FL-based approaches have demonstrated significant potential in critical sectors like smart healthcare, smart transportation, unmanned aerial vehicles (UAVs), smart cities, and intelligent manufacturing. These areas benefit from the enhanced data privacy and distributed computing capabilities enabled by FL, highlighting its role in supporting the complex infrastructures required for next-generation IIoT systems. The practical applications of FL in these fields demonstrate its potential to facilitate more secure, efficient, and scalable solutions for IIoT-driven industrial systems.
IV-D Benchmarks
Benchmarking is crucial for evaluating and improving LLMs. “Continual Instruction Tuning” [231] was proposed as a new evaluation benchmark to assess how machine learning models maintain existing skills while incorporating new knowledge. This benchmark uses ten diverse datasets covering eight task types to test models on their instruction adherence and general knowledge retention, effectively measuring their alignment with human intentions and reasoning abilities.
The SPHINX-X series was developed to enhance the SPHINX framework for increased architectural and training efficiency [252]. This series streamlines the training process, optimizes the architecture by removing unnecessary components, and broadens the framework’s applicability across multiple domains and modalities, including challenging OCR-intensive tasks. This adaptation results in a diverse suite of models that vary in parameter size and support multilingual capabilities.
LAVIS [253] was presented as an open-source deep learning library tailored for language-vision integration. LAVIS provides a unified platform facilitating access to cutting-edge image and video language models, along with a suite of tools for a wide range of multimodal tasks. The library is designed for scalability and flexibility, preparing it for future expansion and customization, with comprehensive benchmarking results demonstrating its effectiveness in language-vision applications.
V Outlook and Future Opportunities
V-A Data Generation
Generative models such as the Diffusion Model [91] have achieved significant success in NLP and computer vision. However, their application in industrial settings remains underexplored. In domains like equipment fault diagnosis and anomaly detection, where small sample sizes are a frequent challenge, the practical implementation and effectiveness of these models require further investigation. The integration of generative models into these specific industrial applications holds potential, but their suitability for handling limited data scenarios is still an open question.
As the demand for AIGC continues to grow, deploying AIGC within mobile edge networks presents a critical area for exploration [7]. These networks offer real-time, personalized AIGC services while ensuring user privacy. However, advancements in supporting technologies and the expansion of practical application scenarios are needed to fully realize the potential of AIGC in these decentralized environments. Addressing these challenges is essential to enable scalable and efficient deployment of AIGC in diverse industrial settings.
Finally, improving instruction adjustment techniques is crucial for enhancing large language models’ ability to interpret and respond accurately to instructions across broader scenarios. Current methods range from manual annotation to leveraging LLMs like the GPT series for automated generation of training data. Although initial results have been promising, integrating these methods into more efficient workflows remains a topic for future research. Furthermore, advancements in generative technologies such as Diffusion Models have increased the realism of synthetic data, including Deepfakes [254]. This creates significant challenges for verifying authenticity, both for human observers and machine classifiers. To mitigate the risks associated with Deepfakes, detection methods are being strengthened, incorporating spatial, frequency, and temporal analyses as well as techniques for identifying synthetic artifacts [255].
V-B Model/Knowledge Reuse
Model reuse employs strategies such as duplication [256], stacking [257] and combining [258] weights of pre-trained smaller models or using knowledge distillation to initialize larger models. The fundamental intuition is that reusing smaller models to warm-start the training of large transformers provides a robust initial point, thereby stabilizing optimization from the outset and conserving data samples needed to explore optimization trajectories. The scaling laws [259] of models can be through model reuse reproduced, even of extreme data conditions scarcity.
Some studies enhance future learning by establishing a knowledge pool [114]. When adapting to new tasks, relevant knowledge is first retrieved from a knowledge base according to the specific distribution of the new task. Subsequently, the pre-trained model is fine-tuned using knowledge transfer methods, such as joint reparameterization, to facilitate effective knowledge reuse. This approach enhances the model’s adaptability to new tasks and optimizes learning efficiency.
V-C Beyond Single Modality
Recent advancements are not confined to unimodal models but have expanded to multimodal pre-trained Models, such as CLIP [20]. These multimodal PTMs are capable of processing, responding to, and reasoning with various types of inputs. While significant progress has been made in visual recognition, particularly in using textual information to select appropriate prompts, interest in going beyond visual recognition is rapidly growing.
For example, to automate the generation of Electrocardiogram reports and ensure their versatility, a Multimodal Electrocardiogram IT framework has been proposed, attempting to utilize LLMs and multimodal instructions to generate Electrocardiogram reports [260]. ImageBind [261] presents a method for joint embedding across six different modalities—images, text, audio, depth, thermal, and IMU data. It extends its zero-shot capabilities to new modalities just by leveraging its natural pairing with images. It enables novel out-of-the-box applications including cross-modal retrieval, composing modalities with arithmetic, cross-modal detection, and generation. These studies mark a significant step forward in multimodal learning, with the integration of multiple modalities opening new possibilities and pathways for future industrial research and applications.
V-D Learning with Restricted Resources
LMs have demonstrated substantial efficacy in industrial applications, yet their continual tuning often demands significant computational resources. The deployment of these models extends beyond cloud-based environments to edge devices, including FPGAs, where models are often quantized and hard-coded to accelerate processing, thus rendering them non-modifiable. Recent advances in CL research, as indicated by Prabhu et al. [262], increasingly focus on applying models in resource-constrained scenarios. This trend is expected to highlight and address key challenges related to computational efficiency in future developments.
Traditional methods typically rely on computationally intensive backpropagation for model updates, which may not be feasible on some devices. As an innovative solution, the Forward-Only Adaptation [263] method adapts models solely at test time using a derivative-free covariance matrix evolution strategy, circumventing the need for backpropagation. This strategy provides a new possibility for model adaptation in environments with limited computational resources.
Edge-based LMs can achieve fast, near real-time analysis and response; however, due to the limited resources of edge nodes and end devices, they struggle to support such large models [34, 264]. Therefore, it is essential to integrate end, edge, and cloud computing to fully leverage their respective advantages, thereby compensating for the resource limitations at any single level.
VI Conclusion
This paper comprehensively surveys and forecasts the development prospects of IIoT-Enhanced General Industrial Intelligence based on large models. Initially, the paper outlines typical application scenarios in the industrial sector and introduces industrial-scale models with significant potential. The discussion is then divided into two main themes: the technological support provided by large models (LMs for GII) for general industrial intelligence and the potential challenges these models face in actual industrial applications (LMs on GII). We explore how pre-trained LMs can be developed to address specific industrial needs and discuss how CL and incremental learning can further enhance the adaptability and sustainability of these models.
Lastly, the paper discusses future research directions, including data generation, model reuse, and multimodal learning. This review aims to inspire more practitioners and researchers to explore continual learning to advance the development of robust and self-evolving industrial AI systems.
VII Acknowledgements
We would like to extend our special thanks to Ruyi Huang, Longya Xiao, Yihao Wang, Chenguang Ma, Xiang Zhang, Qiurui Zhang and Penghong Lu from the Shien-Ming Wu School of Intelligent Engineering, South China University of Technology, China, for their invaluable support in providing data and materials in the field of industrial applications during the writing of this paper.